Grade 12th passMechanics

Two particles A and B make an angle of theta 1 and theta 2 with the line joining A and B find the velocity of approach.

Xyz
7 Years agoGrade 12th pass

1 Answer

Kapil Khare
7 Years ago
Let the velocity of the particle A be u and the velocity of the particle B be v.
So, component of velocity of A along A and B = u*cos\theta1
And, component of velocity of B along A and B = v*cos\theta2
If the direction of these two components are same, then velocity of approach = u*cos\theta1 – v*cos\theta2
If the direction of these two components are same, then velocity of approach = u*cos\theta1 + v*cos\theta2